Circular No. 5993

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1994 No. 2 Yukio Sakurai (Mito, Ibaragi, Japan) reports his discovery of a nova

on SG400 patrol films taken with a 300-mm camera lens and provides the following position: α = 18^{h}27^{m}28^{s}, δ = −17^{o}14′.2 (equinox 1950.0). Magnitude estimates: May 16.733 UT, [11.5 (Sakurai); 20.710, 10.8 (Sakurai); 22.603, 11.4 (S. Takahashi, Shiga, Japan; visual); 22.620, 11.6 (M. Watanabe, Toyama, Japan; visual).

S. J. Austin, G. Schwarz, and S. Starrfield, Arizona State University; and R. M. Wagner, Ohio State University, report: “Optical spectrograms of Nova Sgr 1994 No. 2 were obtained with the Perkins 1.8-m telescope (+ CCD spectrograph) on May 23.3 UT. The spectra (range 473–695 nm, resolution 1.2 nm) exhibit permitted emission lines due to the Balmer series of hydrogen and Fe II (multiplets 37, 38, 42, 48, 49, 55, and 74) superposed on a red continuum. Many of the Fe II lines appear to exhibit weak P-Cyg profiles, but these are much less certain at our resolution. The profile of Hα exhibits an asymmetry to the red. The FWHM of Hα and Hβ emission, corrected for instrumental resolution, was ∼ 450 km/s. The spectra confirm the identification as a nova with an early post-maximum ‘Fe II class’ spectrum. The widths of the emission lines suggest that the development of this nova will be moderately slow.”

MARKARIAN 421 T. Takahashi, M. Tashiro, T. Kamae, and K. Makishima, University

of Tokyo; T. Kii and F. Makino, ISAS; T. Ohashi, F. Takahara, and N. Y. Yamasaki, Tokyo Metropolitan University; R. Hartman and G. Madejski, Goddard Space Flight Center, NASA; R. C. Lamb, Iowa State University; and T. C. Weekes, Whipple Observatory, report: “The BL-Lac object Mkn 421 was observed from the x-ray astronomy satellite ASCA during May 16.4–17.3 UT. The preliminary energy flux obtained from the GIS aboard ASCA was very bright, 1.8 × 10^{−10} erg cm^{−2} s^{−1} (2–10 keV), which is ∼ 2 times higher than that observed by EXOSAT in December 1984 (George et al. 1988, M.N.R.A.S. 232, 793). The photon index was ∼ 2.5. Observations at optical, infrared, radio, and other wavelengths are encouraged.”

COMET MUELLER (1993a) Total visual magnitude estimate by A. Hale (near Orogrande, NM,

0.41-m reflector): May 17.43 UT, 11.3.

1994 May 23 Daniel W. E. Green